# Mirror Descent Policy Optimisation for Robust Constrained Markov Decision Processes

### v1

@@ B.1 Softmax policy gradient with unregularised objective
$V_{\pi_{\theta_{t}},p}(\rho)-V_{\pi_{\theta^{*}},p}(\rho)\leq\frac{2SM_{p}(\pi^{*})^{2}}{U_{p}\eta(1-\gamma)^{3}t}\left\lVert\frac{1}{\mu}\right\rVert_{\infty}\,,$ (59)
where $M_{p}(\pi^{*})=\left\lVert\frac{d_{\mu}^{\pi_{\tau}^{*},p}}{\mu}\right\rVert_{\infty}$.
B.2 Softmax policy gradient with regularised objective
Mei et al., (2020) previously used the above smoothness and continuity to prove the convergence rate for softmax policies on an entropy-regularised objective. Below we summarise their key results that factor into our analysis.
Theorem 6 in Mei et al., (2020) provides convergence rate results for softmax policies with entropy regularisation, a result which can be reused in our setting with a few parameter changes. We rephrase the theorem in slightly more general manner based on the relation between the learning rate $\eta$ and the smoothness coefficient $l_{\theta}$.
Theorem 6 (Convergence rate for softmax policies with entropy regularisation, Theorem 6 in Mei et al., (2020) rephrased).
Let $p$ be the transition dynamics, let $\{\theta_{t}\}_{t\geq 1}$ be a sequence generated by regularised policy gradient over logits according to
$\theta_{t+1}=\theta_{t}-\eta\nabla V_{\pi_{\theta_{t}},p}^{\tau}(\rho)\,,$ (60)
with $\eta=1/l_{\theta}$, let the initial state be sampled from $\mu$, and let $U_{p}=\inf_{s\in\mathcal{S},t\geq 1}\pi_{\theta_{t}}(a^{*}(s|p)|s)>0$. Then, for all $t\geq 1$, the regret on the regularised objective is bounded by
$V_{\pi_{\theta_{t}},p}^{\tau}(\rho)-V_{\pi_{\tau}^{*},p}^{\tau}(\rho)\leq\left\lVert\frac{1}{\mu}\right\rVert_{\infty}\frac{1+\tau\log(A)}{(1-\gamma)^{2}}e^{-C\,(t-1)}\,,$ (61)
where $C=\frac{\eta}{S}\min_{s}\mu(s)U_{p}^{2}(M_{p}(\pi_{\tau}^{*}))^{-1}$ is independent of the iteration $t$.
Appendix C Supporting lemmas for sample-based analysis
C.1 Analysis of the multipliers
@@
The sequence of multipliers produced by Robust PMD-PD $\{\lambda_{k,j}\}_{k\geq 0},\,j\in[m]$ satisfy the following properties:
@@
    Positive modified multiplier: for any macro step $k\geq 0$, $\lambda_{k,j}+\eta_{\lambda}\hat{V}_{\pi_{k},p_{k}}^{j}(\rho)\geq 0$ for all $j\in[m]$.
@@
    Bounded initial multiplier: for macro step 00, $|\lambda_{0,j}|^{2}\leq|\eta_{\lambda}\hat{V}_{\pi_{k},p_{k}}^{j}(\rho)|^{2}$ for all $j\in[m]$
@@
1) Note that $\lambda_{0,j}\geq 0$ is trivially satisfied by the initialisation in Algorithm 1. Via induction, if $\lambda_{k,j}\geq 0$ note that if $\hat{V}_{\pi_{k+1},p_{k+1}}(\rho)\geq 0$ then $\lambda_{k},j+\eta_{\lambda}\hat{V}_{\pi_{k+1},p_{k+1}}(\rho)\geq 0$; if it is negative, then $-\eta_{\lambda}\hat{V}_{\pi_{k+1},p_{k+1}}(\rho)\geq 0$.
@@
Lemma 13 (Lower bound on the inner product, Eq. 41 in Liu et al., (2021)).
@@
$\left\langle\lambda_{k}+\eta_{\lambda}\hat{V}_{\pi_{k},p_{k}}^{1:m}(\rho),\hat{V}_{\pi_{k+1},p_{k+1}}^{1:m}(\rho)\right\rangle$
$\geq\frac{1}{2\eta_{\lambda}}\left(\left\lVert\lambda_{k+1}\right\rVert^{2}-\left\lVert\lambda_{k}\right\rVert^{2}\right)+\frac{\eta}{2}\left(\left\lVert V_{\pi_{k},p_{k}}^{1:m}(\rho)\right\rVert^{2}-\left\lVert V_{\pi_{k+1},p_{k+1}}^{1:m}(\rho)\right\rVert^{2}\right)$
$+\left\langle\lambda_{k},-\epsilon_{k-1}\right\rangle+\left\lVert\epsilon_{k+1}\right\rVert^{2}+\eta_{\lambda}\left\langle\epsilon_{k},V_{\pi_{k},p_{k}}^{i:m}(\rho)\right\rangle-\eta_{\lambda}\left\lVert\epsilon_{k}\right\rVert-2\eta_{\lambda}\left\langle V_{\pi_{k+1},p_{k+1}}^{i:m}(\rho),\epsilon_{k+1}\right\rangle-\frac{\gamma^{2}\eta_{\lambda}}{(1-\gamma)^{4}}B_{d_{\rho}^{\pi_{k+1},p_{k+1}}}(\pi_{k+1},\pi_{k})\,.$
C.2 Analysis of the Bregman divergence
@@
Lemma 14 (Pushback property, Lemma 2 in Liu et al., (2021)).
@@
By selecting $z=\pi$, $y=\pi_{k}^{*}$, and $x^{*}=\pi^{*}$, and using the weighted Bregman divergence, it is possible to use the property in the context of softmax policies, which are always in the interior of the probability simplex.
Lemma 15 (Pushback property for softmax policies, Lemma 10 in Liu et al., (2021)).
For any softmax policy $\pi$ and any $p$,
$\mathbf{V}_{\pi_{k}^{*},p}(\rho)+\frac{\alpha}{1-\gamma}B_{d_{\rho}^{\pi_{k}^{*},p}}(x^{*},y)\leq\mathbf{V}_{\pi,p}(\rho)+\alpha\left(B_{d_{\rho}^{\pi,p}}(\pi,\pi_{k})-B_{d_{\rho}^{\pi,p}}(\pi,\pi_{k}^{*})\right).$
The weighted Bregman divergence under the generating function $\sum_{i}p(i)\log(p(i))$ is equivalent to the KL-divergence, which is bounded by $\log(A)$ as shown below by noting that the uniform policy is the maximum entropy policy.
@@
Let $\pi_{0}$ be a uniform policy, then for any policy $\pi\in\Pi$ and any $p\in\mathcal{P}$,
$B_{d_{\rho}^{\pi,p}}(\pi,\pi_{0})=\sum_{s\in\mathcal{S}}d_{\rho}^{\pi,p}(s)\sum_{a\in\mathcal{A}}\pi(a|s)\log(A\pi(a|s))\leq\log(A)\,.$
C.3 Convergence of approximate entropy-regularised NPG
@@
Lemma 18 (Number of inner loop iterations, Lemma 7 in Liu et al., (2021)).
Let $\epsilon>0$, $\left\lVert\hat{\mathbf{Q}}_{\pi_{k}^{t},p}^{\alpha}-\mathbf{Q}_{\pi_{k}^{t},p}^{\alpha}\right\rVert_{\infty}\leq\epsilon$, $\eta\leq(1-\gamma)/\alpha$, $t_{k}=\frac{1}{\eta\alpha}\log(3C_{k}K)$, and
@@
$\left\lVert\mathbf{V}_{\pi_{k}^{*},p}^{\alpha}-\mathbf{V}_{\pi_{k+1}}{\alpha}\right\rVert_{\infty}\leq\frac{1}{K}+\frac{6\epsilon}{(1-\gamma)^{2}}$ (67)
@@
Lemma 19 (Performance difference lemma, Lemma 4 in Cen et al., (2022) and Lemma 6 in Liu et al., (2021)).
@@
$\mathbf{V}_{\pi_{k}^{t+1},p}^{\alpha}(\rho)-\mathbf{V}_{\pi_{k}^{t},p}^{\alpha}(\rho)\leq\frac{2}{1-\gamma}\left\lVert\hat{\mathbf{Q}}_{\pi_{k}^{t},p}^{\alpha}-\mathbf{Q}_{\pi_{k}^{t},p}^{\alpha}\right\rVert_{\infty}\,.$ (69)
Corollary 1.
It follows from Lemma 19 that
$Q_{\pi_{k}^{t+1},p}(s,a)-Q_{\pi_{k}^{t},p}(s,a)=\gamma(V_{\pi_{k}^{t+1},p}(p(\cdot|s,a))-V_{\pi_{k}^{t},p}(p(\cdot|s,a)))\leq\frac{2\gamma}{1-\gamma}\left\lVert\hat{Q}_{\pi_{k}^{t}}-Q_{\pi_{k}^{t}}\right\rVert_{\infty}\,.$ (70)
C.4 Performance difference across transition kernels

### v2

@@ C.2 Regret bound for approximate TMA
$\hskip 113.81102pt\frac{1}{1-\gamma}\sum_{s\in\mathcal{S}}d_{\rho}^{\pi,p^{t+1}}(s)\sum_{a\in\mathcal{A}}\pi(a|s)\langle p^{t}(\cdot|s,a)-p^{t+1}(\cdot|s,a),\hat{G}_{\pi,p^{t}}(s,a,\cdot)-G_{\pi,p^{t}}(s,a,\cdot)\rangle\Big)$ (definition of $M=\sup_{p\in\mathcal{P}}\left\lVert\frac{d_{\rho}^{\pi,p^{*}}}{d_{\rho}^{\pi,p}}\right\rVert_{\infty}$ and again Eq. 28 )
$=M\Big(\left(V_{\pi,p^{t}}(\rho)-V_{\pi,p^{t+1}}(\rho)\right)+$
$\hskip 113.81102pt\frac{1}{1-\gamma}\sum_{s\in\mathcal{S}}d_{\rho}^{\pi,p^{t+1}}(s)\sum_{a\in\mathcal{A}}\pi(a|s)\langle p^{t}(\cdot|s,a)-p^{t+1}(\cdot|s,a),\hat{G}_{\pi,p^{t}}(s,a,\cdot)-G_{\pi,p^{t}}(s,a,\cdot)\rangle\Big)$ (first performance difference lemma (Lemma 22))
$\geq M\left(V_{\pi,p^{t}}(\rho)-V_{\pi,p^{t+1}}(\rho)-\frac{2\epsilon^{\prime}}{1-\gamma}\right)\,$,
where the last step follows from a bound of the absolute value as in the proof of Lemma 5.
Combining both A) and B), denoting $\delta_{t}=V_{\pi,p^{*}}(\rho)-V_{\pi,p^{t}}(\rho)$, and rearranging, we obtain
$\delta_{t}+M\left(\delta_{t+1}-\delta_{t}-\frac{2\epsilon^{\prime}}{1-\gamma}\right)\leq\frac{\alpha_{p}(t)}{\eta_{p}(t)(1-\gamma)}\left(B_{d_{\rho}^{\pi,p^{*}}}(p^{*},p^{t})-B_{d_{\rho}^{\pi,p^{*}}}(p^{*},p^{t+1})\right)+\frac{2\epsilon^{\prime}}{1-\gamma}$
$M\delta_{t+1}+(1-M)\delta_{t}\leq\frac{\alpha_{p}(t)}{\eta_{p}(t)(1-\gamma)}\left(B_{d_{\rho}^{\pi,p^{*}}}(p^{*},p^{t})-B_{d_{\rho}^{\pi,p^{*}}}(p^{*},p^{t+1})\right)+\frac{2\epsilon^{\prime}(1+M)}{1-\gamma}$ (rearrange)
$\left(\frac{1}{M}-1\right)\delta_{t}+\delta_{t+1}\leq\frac{\alpha_{p}(t)}{M\eta_{p}(t)(1-\gamma)}\left(B_{d_{\rho}^{\pi,p^{*}}}(p^{*},p^{t})-B_{d_{\rho}^{\pi,p^{*}}}(p^{*},p^{t+1})\right)+\frac{4\epsilon^{\prime}}{1-\gamma}$ (divide by $M$ and note $\frac{1}{M}+1<2$)
$\delta_{t+1}+\frac{\alpha_{p}(t)}{M\eta_{p}(t)(1-\gamma)}B_{d_{\rho}^{\pi,p^{*}}}(p^{*},p^{t+1})\leq\left(1-\frac{1}{M}\right)\left(\delta_{t}+\frac{\alpha_{p}(t)}{(M-1)\eta_{p}(t)(1-\gamma)}B_{d_{\rho}^{\pi,p^{*}}}(p^{*},p^{t})\right)+\frac{4\epsilon^{\prime}}{1-\gamma}$ (rearrange)
$\delta_{t+1}+\frac{\alpha_{p}(t+1)}{(M-1)\eta_{p}(t+1)(1-\gamma)}B_{d_{\rho}^{\pi,p^{*}}}(p^{*},p^{t+1})\leq\left(1-\frac{1}{M}\right)\left(\delta_{t}+\frac{\alpha_{p}(t)}{(M-1)\eta_{p}(t)(1-\gamma)}B_{d_{\rho}^{\pi,p^{*}}}(p^{*},p^{t})\right)+\frac{4\epsilon^{\prime}}{1-\gamma}\,.$ (since $M\geq\frac{1}{1-\gamma}$, it follows that $\eta_{p}(t+1)\geq\eta_{p}(t)/\gamma\geq\eta_{p}(t)\frac{M}{M-1}$)
Using Lemma 11 with $\beta=1-\frac{1}{M}$, $x_{t}=\delta_{t}+\frac{\alpha_{p}(t)}{(M-1)\eta_{p}(t)(1-\gamma)}B_{d_{\rho}^{\pi,p^{*}}}(p^{*},p^{t})$, and $b=\frac{4\epsilon^{\prime}}{1-\gamma}$, it follows that
$\delta_{t}$ $\leq\delta_{t}+\frac{\alpha_{p}(t)}{(M-1)\eta_{p}(t)(1-\gamma)}B_{d_{\rho}^{\pi,p^{*}}}(p^{*},p^{t})$
$\leq\left(1-\frac{1}{M}\right)^{t}\left(\delta_{0}+\frac{\alpha_{p}(0)}{(M-1)\eta_{p}(0)(1-\gamma)}B_{d_{\rho}^{\pi,p^{*}}}(p^{*},p^{0})\right)+\frac{4M\epsilon^{\prime}}{1-\gamma}$
$\leq\left(1-\frac{1}{M}\right)^{t}\frac{3C}{1-\gamma}+\frac{4M}{1-\gamma}\epsilon^{\prime}\,$,
where the last step follows from $\delta_{0}\leq\frac{2C}{1-\gamma}$ and $\eta_{p}(0)\geq\alpha_{p}(0)\frac{\gamma}{C(1-\gamma)}B_{d_{\rho}^{\pi,p^{0}}}(p^{*},p^{0})$. ∎
C.3 Approximate action next-state value
The proof of Lemma 6 is given below.
Proof.
By triangle inequality, the error $\left\lVert\hat{G}-G\right\rVert_{\infty}$ is upper bounded by the truncation error and the estimation error, so it suffices to bound both by half of the desired error in Eq. 32.
For the truncation error, observe that
$|\mathbb{E}\left[\gamma^{t}c(s,a,s^{\prime})\right]|\leq\sum_{t=N_{G,t}}^{\infty}C\gamma^{t}=C\frac{\gamma^{N_{G,t}}}{1-\gamma}\,.$
For the estimation error of any particular $(s,a,s^{\prime})\in\mathcal{S}\times\mathcal{A}\times\mathcal{S}$ and any $t\in\{0,\dots,t_{k}^{\prime}\}$, apply Hoeffding’s inequality with the random variable $X_{t}(s,a,s^{\prime})=\sum_{l=0}^{N_{G,t}-1}\gamma^{l}\left[c(s_{l}^{j},a_{l}^{j})\right]\in[-\frac{C}{1-\gamma},\frac{C}{1-\gamma}]$ conditioned on $(s_{0},a_{0},s_{1})=(s,a,s^{\prime})$, error tolerance $\zeta=\frac{\gamma^{N_{G,t}}C}{1-\gamma}$, and failure probability $\delta^{\prime}=\frac{\delta}{S^{2}At_{k}^{\prime}}$. This results in
$P(|\bar{X}(s,a,s^{\prime})-\mathbb{E}[\bar{X}(s,a,s^{\prime})]|\geq\zeta)\leq 2\exp\left(-\frac{M_{G,t}\zeta^{2}(1-\gamma)^{2}}{2C^{2}}\right)=\delta^{\prime}$
$M_{G,t}\leq\frac{2C^{2}}{\zeta^{2}(1-\gamma)^{2}}\log\left(\frac{2}{\delta^{\prime}}\right)=\frac{2\gamma^{2N_{G,t}}}{(1-\gamma)^{2}}\log\left(\frac{2}{\delta^{\prime}}\right)\,.$
Repeating the same procedure for all $\mathcal{S}\times\mathcal{A}\times\mathcal{S}$, and applying union bound, we obtain a failure probability
$P\left(\bigcup_{s,a,s^{\prime},t}|\bar{X}_{t}(s,a,s^{\prime})-\mathbb{E}[\bar{X}_{t}(s,a,s^{\prime})]|\geq\zeta\right)$
$\leq\sum_{s,a,s^{\prime},t}P(|\bar{X}_{t}(s,a,s^{\prime})-\mathbb{E}[\bar{X}_{t}(s,a,s^{\prime})]|\geq\zeta)$
$\leq\sum_{s,a,s^{\prime},t}\delta^{\prime}=\delta\,.$
This concludes the proof that the setting $M_{G,t}=\frac{2\gamma^{-2N_{G,t}}}{(1-\gamma)^{2}}\log\left(\frac{2}{\delta^{\prime}}\right)$ is sufficient to obtain an estimation error of at most $C\frac{\gamma^{N_{G,t}}}{1-\gamma}$ with a failure probability of at most $\delta$ across all iterations and state-action-state triplets. ∎
C.4 Sample complexity of Approximate TMA
The proof of Lemma 7 is given below.
Proof.
Note that it suffices to bound both parts in the RHS of Eq. 31 with $\epsilon^{\prime}/2$.
For the first part of Eq. 31, note that requiring
$\left(1-\frac{1}{M}\right)^{t}\frac{3C}{1-\gamma}\leq\frac{\epsilon^{\prime}}{2}$ (50)
implies that the number of iterations is bounded by
$t\leq\log\left(\frac{\epsilon^{\prime}(1-\gamma)}{6C}\right)/\log\left(\frac{M-1}{M}\right)=\log_{\frac{M}{M-1}}\left(\frac{6C}{\epsilon^{\prime}(1-\gamma)}\right)\,.$
This confirms that the setting $t_{k}^{\prime}\geq\log_{\frac{M}{M-1}}\left(\frac{6C}{\epsilon^{\prime}(1-\gamma)}\right)$ is sufficient to satisfy Eq. 50.
For the second part of Eq. 31, define $\epsilon^{\prime}=2C\frac{\gamma^{H}}{1-\gamma}$ consistent with Eq. 29. This amounts to requiring that
$\frac{4M}{1-\gamma}\epsilon^{\prime}=\frac{8MC}{(1-\gamma)^{2}}\gamma^{H}\leq\epsilon^{\prime}/2$ (51)
This implies a requirement of at most
$H\leq\log\left(\frac{\epsilon^{\prime}(1-\gamma)^{2}}{16MC}\right)/\log(\gamma)=\log_{\frac{1}{\gamma}}\left(\frac{16MC}{\epsilon^{\prime}(1-\gamma)^{2}}\right)\,.$
Therefore the setting of $H\geq\log_{\frac{1}{\gamma}}\left(\frac{16MC}{\epsilon^{\prime}(1-\gamma)^{2}}\right)$ suffices to obtain Eq. 51.
With the settings of $M_{G}$, $t_{k}^{\prime}$, and $H$, note that $\gamma^{-2H}=\left(\frac{1}{\gamma}\right)^{2\log_{\frac{1}{\gamma}}\left(\frac{16MC}{\epsilon^{\prime}(1-\gamma)^{2}}\right)}=\left(\frac{16MC}{\epsilon^{\prime}(1-\gamma)^{2}}\right)^{2}$. Therefore
$M_{G}$ $=\frac{2\gamma^{-2H}}{(1-\gamma)^{2}}\log\left(\frac{2S^{2}At_{k}^{\prime}}{\delta}\right)$
$=2\left(\frac{16M}{(1-\gamma)^{2}\epsilon^{\prime}}\right)^{2}\log\left(\frac{2S^{2}At_{k}^{\prime}}{\delta}\right)\,.$
The total number of samples is determined by the number of state-action-state triplets, the number of iterations, the horizon, and the number of trajectories per mini-batch:
$n$ $=|S|^{2}\,|A|\,t_{k}^{\prime}HM_{G}$
$\leq S^{2}A\log_{\frac{M}{M-1}}\left(\frac{6C}{\epsilon^{\prime}(1-\gamma)}\right)\log_{\frac{1}{\gamma}}\left(\frac{16M}{(1-\gamma)^{2}\epsilon^{\prime}}\right)2\left(\frac{16M}{(1-\gamma)^{2}\epsilon^{\prime}}\right)^{2}\log\left(\frac{2t_{k}^{\prime}S^{2}A}{\delta}\right)$
$=512S^{2}A\frac{M^{2}}{(1-\gamma)^{4}\epsilon^{\prime\prime 2}}\log_{\frac{M}{M-1}}\left(\frac{6C}{\epsilon^{\prime}(1-\gamma)}\right)\log\left(\frac{16M}{(1-\gamma)^{4}\epsilon^{\prime}}\right)\log\left(\frac{2t_{k}^{\prime}S^{2}A}{\delta}\right)$
$=\tilde{\mathcal{O}}\left(\frac{S^{2}AM^{2}}{(1-\gamma)^{4}\epsilon^{\prime 2}}\right)\,.$
Appendix D Supporting lemmas and proofs for Robust Sample-based PMD-PD
D.1 Analysis of the multipliers
@@
The sequence of multipliers produced by Robust Sample-based PMD-PD $\{\lambda_{k,j}\}_{k\geq 0},\,j\in[m]$ satisfy the following properties:
@@
    Positive augmented multiplier: for any macro step $k\geq 0$, $\lambda_{k,j}+\eta_{\lambda}\hat{V}_{\pi_{k},p_{k}}^{j}(\rho)\geq 0$ for all $j\in[m]$.
@@
    Bounded initial multiplier: for macro step 0, $|\lambda_{0,j}|^{2}\leq|\eta_{\lambda}\hat{V}_{\pi_{0},p_{0}}^{j}(\rho)|^{2}$ for all $j\in[m]$
@@
1) Note that $\lambda_{0,j}\geq 0$ is trivially satisfied by the initialisation in Algorithm 2. Via induction, if $\lambda_{k,j}\geq 0$ note that if $\hat{V}_{\pi_{k+1},p_{k+1}}(\rho)\geq 0$ then $\lambda_{k},j+\eta_{\lambda}\hat{V}_{\pi_{k+1},p_{k+1}}(\rho)\geq 0$; if it is negative, then $\eta_{\lambda}\hat{V}_{\pi_{k+1},p_{k+1}}(\rho)\geq 0$.
@@
Lemma 14 (Lower bound on the inner product, Eq. 41 in Liu et al., 2021b ).
@@
$\left\langle\lambda_{k}+\eta_{\lambda}\hat{V}_{\pi_{k},p_{k}}^{1:m}(\rho),V_{\pi_{k+1},p_{k+1}}^{1:m}(\rho)\right\rangle$
$\geq\frac{1}{2\eta_{\lambda}}\left(\left\lVert\lambda_{k+1}\right\rVert^{2}-\left\lVert\lambda_{k}\right\rVert^{2}\right)+\frac{\eta_{\lambda}}{2}\left(\left\lVert V_{\pi_{k},p_{k}}^{1:m}(\rho)\right\rVert^{2}-\left\lVert V_{\pi_{k+1},p_{k+1}}^{1:m}(\rho)\right\rVert^{2}\right)$
$+\left\langle\lambda_{k},-\epsilon_{k-1}\right\rangle+\left\lVert\epsilon_{k+1}\right\rVert^{2}+\eta_{\lambda}\left\langle\epsilon_{k},V_{\pi_{k},p_{k}}^{1:m}(\rho)\right\rangle-\eta_{\lambda}\left\lVert\epsilon_{k}\right\rVert-2\eta_{\lambda}\left\langle V_{\pi_{k+1},p_{k+1}}^{1:m}(\rho),\epsilon_{k+1}\right\rangle-\frac{\gamma^{2}\eta_{\lambda}}{(1-\gamma)^{4}}B_{d_{\rho}^{\pi_{k+1},p_{k+1}}}(\pi_{k+1},\pi_{k})\,.$
D.2 Analysis of the Bregman divergence
@@
Lemma 15 (Pushback property, Lemma 2 in Liu et al., 2021b ).
@@
By selecting $z=\pi$, $y=\pi_{k}^{*}$, and $x^{*}=\pi^{*}$, and using the weighted Bregman divergence, it is possible to use the property in the context of policies in the interior of the probability simplex.
Lemma 16 (Pushback property, Lemma 11 in Liu et al., 2021b rephrased).
For any policy $\pi\in\text{Int}(\Delta(\mathcal{A}))$ and any $p$,
$\mathbf{V}_{\pi_{k}^{*},p}(\rho)+\frac{\alpha}{1-\gamma}B_{d_{\rho}^{\pi_{k}^{*},p}}(\pi_{k}^{*},\pi)\leq\mathbf{V}_{\pi,p}(\rho)+\alpha\left(B_{d_{\rho}^{\pi,p}}(\pi,\pi_{k})-B_{d_{\rho}^{\pi,p}}(\pi,\pi_{k}^{*})\right).$
The weighted Bregman divergence under the generating function $\sum_{i}p(i)\log(p(i))$ is equivalent to the KL-divergence, which is bounded by $\log(A)$ if the reference policy is uniform.
@@
Let $\pi^{\prime}$ be a uniform policy, then for any policy $\pi\in\Pi$ and any $p\in\mathcal{P}$,
$B_{d_{\rho}^{\pi,p}}(\pi,\pi^{\prime})=\sum_{s\in\mathcal{S}}d_{\rho}^{\pi,p}(s)\sum_{a\in\mathcal{A}}\pi(a|s)\log(A\pi(a|s))\leq\log(A)\,.$
D.3 Convergence of approximate entropy-regularised NPG
@@
Lemma 19 (Number of inner loop iterations, Lemma 7 in Liu et al., 2021b ).
Let $\epsilon>0$, $\left\lVert\hat{\mathbf{Q}}_{\pi_{k}^{t},p_{k}}^{\alpha}-\mathbf{Q}_{\pi_{k}^{t},p_{k}}^{\alpha}\right\rVert_{\infty}\leq\epsilon$, $\eta\leq(1-\gamma)/\alpha$, $t_{k}=\frac{1}{\eta\alpha}\log(3C_{k}K)$, and
@@
$\left\lVert\mathbf{V}_{\pi_{k}^{*},p_{k}}^{\alpha}-\mathbf{V}_{\pi_{k+1},p_{k}}{\alpha}\right\rVert_{\infty}\leq\frac{1}{K}+\frac{6\epsilon}{(1-\gamma)^{2}}$ (57)
@@
Lemma 20 (Performance difference lemma, Lemma 4 in Cen et al., (2022) and Lemma 6 in Liu et al., 2021b ).
@@
$\mathbf{V}_{\pi_{k}^{t+1},p_{k}}^{\alpha}(\rho)-\mathbf{V}_{\pi_{k}^{t},p_{k}}^{\alpha}(\rho)\leq\frac{2}{1-\gamma}\left\lVert\hat{\mathbf{Q}}_{\pi_{k}^{t},p_{k}}^{\alpha}-\mathbf{Q}_{\pi_{k}^{t},p_{k}}^{\alpha}\right\rVert_{\infty}\,.$ (59)
D.4 Performance difference across transition kernels